# Cognitive Hierarchy in Day-to-day Network Flow Dynamics

### v1

@@ 6 CH-Logit Dynamic
If the day-to-day operator $H[\cdot]$ is replaced by a “stochastic” Logit dynamic (Fudenberg and Levine, 1998; Sandholm, 2010; Xiao et al., 2019), we have the following CH-Logit dynamic:
@@
where $\Phi^{\theta}\left[\boldsymbol{z}\right]\equiv\left(d_{w}\Phi^{\theta}_{w}\left[\boldsymbol{z}_{w}\right],w\in W\right)^{T}$, $\Phi^{\theta}_{w}\left[\boldsymbol{z}_{w}\right]\equiv(d_{w}\varphi_{rw}(\boldsymbol{z}_{w}),r\in R_{w})^{T}$, $\varphi_{rw}(\boldsymbol{z}_{w})=\frac{\exp(-\theta z_{rw})}{\sum_{s\in R_{w}}{\exp(-\theta z_{sw})}}$, and $\theta$ is a dispersal parameter capturing travelers’ perception errors. For each $w\in W$, as $\theta\to 0$, choices become equiprobable among all the routes in $R_{w}$ and as $\theta\to\infty$, choices become extremely concentrated on the least cost route of $R_{w}$. Note that for the Logit dynamic, the flow variable (i.e., the first variable) in $\boldsymbol{y}[\cdot]$ of (1) is not explicitly required.
Based on the CH modeling idea in Section 2.3, the predicted flow pattern of each class $k\in\{0,1,2\}$ is given by:
@@
A significantly large $\hat{\theta}$ can be interpreted as a scenario where higher-step travelers anticipate that lower-step travelers will strongly favor yesterday’s shortest path. This situation closely resembles the case of a large $\hat{\gamma}$ in the CH-NTP model. In both cases, higher-step travelers are making an extreme prediction about the responsiveness of lower-step travelers to cost differences.
@@ 6.1 Mixed prediction-based stochastic equilibria
Similar to the case of the CH-NTP dynamic, we term fixed points of the CH-Logit dynamic (25)-(28) as mixed prediction-based stochastic equilibria (MPSE) and describe them in the following proposition.
@@
When the route cost function $\boldsymbol{c}\left(\boldsymbol{x}\right)$ is continuous, the dynamical system (25)-(28) admits at least one MPSE (i.e., one fixed point). Moreover, a vector $\boldsymbol{x}^{\circ}\equiv(\boldsymbol{x}^{k,\circ},k=0,1,2)^{T}$ is an MPSE if $\forall r\in R_{w},w\in W,k\in K$,
@@
where $\boldsymbol{\pi}^{k,\circ}\equiv(\pi_{sw}^{k},s\in R_{w},w\in W)^{T}$ is a function of the aggregate flow pattern $\tilde{\boldsymbol{x}}^{\circ}=\sum_{k}{\boldsymbol{x}^{k,\circ}}$, dictated by (26)-(28).
@@
The existence of a fixed point is guaranteed by the continuities of the closed-form Logit operator and the route cost function. Setting the RHS of (25) to be 0 and rearrange the results yields (29). ∎
@@
$x_{rw}^{k,\star}=p^{k}d_{w}\frac{\exp\left(-\theta c_{rw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}{\sum_{s\in R_{w}}{\exp\left(-\theta c_{sw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}},$ (30)
where $\tilde{\boldsymbol{x}}^{\star}=\sum_{k}{\boldsymbol{x}^{k,\star}}$.
@@
When $\hat{\theta}=\theta$, a $|K|(=3)$-class SUE in Definition 2 parameterized by $\theta$ is an MPSE (i.e., a fixed point) of the dynamical system (25)-(28) , but not vice versa.
@@
At the SUE, first note from (26) that $\boldsymbol{\pi}^{0,\star}=\tilde{\boldsymbol{x}}^{\star}$. Thus, for each $r\in R_{w},w\in W$, a 1-step traveler’s predicted flow (27) is:
$\pi^{1,\star}_{rw}=\hat{\alpha}q_{1}^{0}d_{w}\frac{\exp\left(-\theta c_{rw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}{\sum_{s\in R_{w}}{\exp\left(-\theta c_{sw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}}+(1-\hat{\alpha})q_{1}^{0}\tilde{x}_{rw}^{\star}=\hat{\alpha}\tilde{x}_{rw}^{\star}+(1-\hat{\alpha})\tilde{x}_{rw}^{\star}=\tilde{x}_{rw}^{\star}.$ (31)
where the second equality is due to Definition 2. Accordingly, for each $r\in R_{w},w\in W$, a 2-step traveler’s predicted flow (28) at the SUE is:
@@
Substituting the above predicted costs, $\tilde{\boldsymbol{x}}^{\star}$, into (25) and by Definition 2, (25) becomes 0. Hence, the SUE is an MPSE.
Similar to the CH-NTP dynamic, there exist some MPSE that are not SUE when $|K|\geq 2$. We can easily construct a counter-example demonstrating their existence (similar to Figure 5(a)) to negate the necessity. This example is omitted to conserve space. ∎
@@ 6.2 Local stability
Due to the existence of multiple equilibria, this section investigates the CH-Logit dynamic’s local stability. We first give the Jacobian of the CH-Logit dynamic at the MPSE and then analyze the stability near the SUE.
@@ 6.2.1 Jacobian matrix
The result for a single $w$ can be obtained from Gao and Pavel (2017) after flipping the positive sign due to the negative utility of travel cost. The independence of routes from different OD pairs makes the non-diagonal blocks zero. $\Upsilon^{\theta}[\boldsymbol{c}]$ is NSD because a block diagonal matrix is NSD if and only if each diagonal block is NSD. ∎
@@
The derivation can be found in Appendix G. ∎
@@ 6.2.2 Stability near the SUE
If $\theta\neq\hat{\theta}$, the block submatrices in (34) would not commute with each other, and thus $|\lambda{I}-J\Phi|$ entails directly calculating the inverse matrix of either one of these submatrices. However, none of these inverse submatrices has a closed-form. Hence, we analytically examine the case of $\theta=\hat{\theta}$ only.
@@
Denote $\rho_{i}$ as the $i$-th eigenvalue of $\Upsilon^{\theta}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\star}\right)\right]D\left[\tilde{\boldsymbol{x}}^{\star}\right]$ at the SUE under Assumption 1 and $\rho_{i}\leq 0,\forall i$. Define function $\psi(\rho_{i};\alpha,\allowbreak\hat{\alpha},p^{0},p^{1},p^{2})\equiv\alpha\hat{\alpha}^{2}p^{2}\frac{p^{1}}{p^{0}+p^{1}}\rho_{i}^{3}+(\alpha\hat{\alpha}-\alpha\hat{\alpha}p^{0}-\alpha\hat{\alpha}^{2}p^{2}\frac{p^{1}}{p^{0}+p^{1}})\rho_{i}^{2}+(\alpha\hat{\alpha}p^{0}-\alpha\hat{\alpha}+\alpha)\rho_{i}+(1-\alpha)$ where $p^{0}+p^{1}+p^{2}=1$. When $\hat{\theta}=\theta$ and $|K|\leq 3$, the SUE is locally asymptotically stable when $-1<\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})<1,\forall i$. If $|K|=2$, then $p^{2}=0$. If $|K|=1$, then $p^{1}=p^{2}=0$, $p^{0}=1$ and $\hat{\alpha}=\alpha$.
@@ 6.2.2.1   Perfect prediction
Figure 11: Stable and unstable regions of the CH-Logit dynamic with $\hat{\alpha}=\alpha$ w.r.t. different eigenvalues $\rho_{i}$.
@@
    When $\rho_{i}\geq-1$, the system is always stable irrespective of any parameter. To explain, note that $\psi(\rho_{i}=-1;\alpha,\hat{\alpha}=\alpha,p^{0},p^{1},p^{2})=2\alpha^{2}-2(p^{0}+1)\alpha+1=(\sqrt{2}\alpha-\frac{\sqrt{2}}{2}(1+p^{0}))^{2}-\frac{1}{2}(1+p^{0})^{2}+1$, which is greater than $-1$ because $p^{0}\leq 1$ and $\alpha<1$. Further note that when $\rho_{i}\in[-1,0]$, the function monotonically increases in $\rho_{i}$. Hence, if $\rho_{i}\geq-1$, the cubic function is always greater than -1 but less than $1-\alpha$;
@@
    Given a $\rho_{i}$ in $(-\infty,-1)$, the stable region would gradually shrink as $\alpha$ increases. Moreover, the stable region would be separated into two pieces when $\alpha$ exceeds a certain threshold (approximately 0.6);
@@
For case (i) of $|K|=2$ in Proposition 8, solving $11-\alpha+\frac{\alpha(1-\hat{\alpha}+\hat{\alpha}p^{0})^{2}}{4\hat{\alpha}(p^{0}-1)}>-$ 1 with $\hat{\alpha}\geq\alpha$ yields $p^{0}\leq h(\alpha,\hat{\alpha})\equiv 2\sqrt{2}\sqrt{-\frac{\alpha-2}{\alpha^{2}\hat{\alpha}^{2}}}+\frac{\alpha\hat{\alpha}+\alpha-4}{\alpha\hat{\alpha}}$. Noting that $h(\alpha,\hat{\alpha})$ is monotonically increasing in $\hat{\alpha}$ and $h(\alpha,\hat{\alpha}=\alpha)\geq 2(\sqrt{2}-1)$, we obtain $h(\alpha,\hat{\alpha})\geq 2(\sqrt{2}-1)\approx 0.828$. The stability condition for $|K|=1$ requires all the $\rho_{i}>\frac{\alpha-2}{\alpha}$ while the condition for $|K|=2$ needs all the $\rho_{i}>\frac{1}{(p^{0}-1)\hat{\alpha}}$. Comparing these two thresholds yields: $\frac{1}{(p^{0}-1)\hat{\alpha}}-\frac{\alpha-2}{\alpha}<0$ when $p^{0}>g(\alpha,\hat{\alpha})\equiv\frac{\alpha\hat{\alpha}+\alpha-2\hat{\alpha}}{\alpha\hat{\alpha}-2\hat{\alpha}}$ and $\frac{1}{(p^{0}-1)\hat{\alpha}}-\frac{\alpha-2}{\alpha}>0$ when $p^{0}<g(\alpha,\hat{\alpha})$. Hence, the stable region is expanded when $g(\alpha,\hat{\alpha})<p^{0}<h(\alpha,\hat{\alpha})$, and shrunk when $0<p^{0}<g(\alpha,\hat{\alpha})$.
For case (ii) of $|K|=2$, $p^{0}>h(\alpha,\hat{\alpha})$. Combined size of the two separated stable regions over $\rho_{i}$ is $f_{0}(\alpha,\hat{\alpha},p^{0})-\frac{1}{(p-1)\hat{\alpha}}+0-f_{1}(\alpha,\hat{\alpha},p^{0})=\frac{1}{\hat{\alpha}(1-p^{0})}\cdot\left(1-\sqrt{\frac{8\hat{\alpha}(p^{0}-1)+\alpha(1+\hat{\alpha}-p^{0}\hat{\alpha})^{2}}{\alpha}}\right)$, which is always less than $\frac{2-\alpha}{\alpha}$ when $p^{0}>h(\alpha,\hat{\alpha})$. Hence, the total size of the two separated stable regions is expanded.
Figure 12 summarizes the above discussion by depicting how the stable region alters when $|K|$ changes from 1 to 2 under different $p^{0}$, $\alpha$, $\hat{\alpha}$. Note how the parametric space of $p^{0}$ and $\alpha$ with a shrinking stable region enlarges as over-prediction becomes more severe (i.e., $\frac{\hat{\alpha}}{\alpha}$ becomes larger).
Figure 12: Sketch of how the stable region alters when $|K|$ changes from 1 to 2 under different $p^{0}$, $\alpha$, $\hat{\alpha}$, and $\hat{\alpha}\geq\alpha$.
@@
  • when $\hat{\alpha}\geq\frac{\alpha}{2-\alpha}$, solving $\frac{1}{(p^{0}-1)\hat{\alpha}}<\frac{\alpha-2}{\alpha}$ yields $p^{0}>g(\alpha,\hat{\alpha})$ and solving $\frac{1}{(p^{0}-1)\hat{\alpha}}>\frac{\alpha-2}{\alpha}$ obtains $p^{0}<g(\alpha,\hat{\alpha})$. Hence, the region is expanded when $g(\alpha,\hat{\alpha})<p^{0}<h(\alpha,\hat{\alpha})$ and shrunk when $p^{0}<g(\alpha,\hat{\alpha})$. It can be verified that $g(\alpha,\hat{\alpha})$ is monotonically increasing in $\hat{\alpha}$ and that $h(\alpha,\hat{\alpha})-g(\alpha,\hat{\alpha})$ is monotonically decreasing in $\hat{\alpha}$ when $10<\alpha,\hat{\alpha}<$ 1. Hence, the parametric space of $p^{0}$, $\alpha$, and $\hat{\alpha}$ ($<\alpha$) rendering the stable region shrunk is smaller than the perfect prediction case, while the parametric space leading to an expanded stable region is larger than the perfect prediction case;
  • when $\hat{\alpha}<\frac{\alpha}{2-\alpha}$, $\frac{1}{(p^{0}-1)\hat{\alpha}}<\frac{\alpha-2}{\alpha}$ always holds and thus the expanded region is simply $p^{0}<h(\alpha,\hat{\alpha})$. When prediction behaviors are considered, no stable region of $|K|=1$ is shrunk.
For case (ii) where $f^{\ast}(\alpha,\hat{\alpha},p^{0})<-1$, $p^{0}>h(\alpha,\hat{\alpha})$, it is verified that $f_{0}(\alpha,\hat{\alpha},p^{0})-\frac{1}{(p-1)\hat{\alpha}}+0-f_{1}(\alpha,\hat{\alpha},p^{0})>\frac{2-\alpha}{\alpha}$. Hence, when $p^{0}>h(\alpha,\hat{\alpha})$, the stable region is separated and the combined stable region size increased. Since $h(\alpha,\hat{\alpha})$ is monotonically increasing in $\hat{\alpha}$, the parametric space rendering the total size of the stable region greater is larger than the perfect prediction case.
@@
Figure 13: Sketch of how the stable region changes when $|K|$ changes from 1 to 2 under different $p^{0}$, $\alpha$, $\hat{\alpha}$ and $3-2\sqrt{2}<\hat{\alpha}\leq\alpha$.
@@
This scenario depends on the comparison between $\frac{8\hat{\alpha}}{(1+\hat{\alpha})^{2}}$ and $\alpha$. When $\frac{8\hat{\alpha}}{(1+\hat{\alpha})^{2}}>\alpha$, discussion is the same as the case of $\hat{\alpha}>3-2\sqrt{2}$; i.e., it helps stabilize the dynamic. When $\frac{8\hat{\alpha}}{(1+\hat{\alpha})^{2}}<\alpha$, $f^{\ast}(\alpha,\hat{\alpha},p^{0})<-1$ always holds and thus the stable region is separated, while the combined size is increased. The latter case is similar to the CH-NTP dynamic with $\hat{\gamma}<\frac{\gamma}{2}$.
Summary
The CH-Logit dynamic bears similarity to the CH-NTP dynamic on how under- and over-predictions affect the stability near the UE. Over-predictions ($\hat{\alpha}>\alpha$ for CH-Logit dynamic and $\hat{\gamma}>\gamma$ for CH-NTP dynamic) render the system less stable while mild under-predictions ($3-2\sqrt{2}<\hat{\alpha}<\hat{\alpha}$ for CH-Logit dynamic and $\frac{\gamma}{2}<\hat{\gamma}<\gamma$ for CH-NTP dynamic) help stabilize the small perturbations near the UE. Severe under-predictions ($\hat{\alpha}<3-2\sqrt{2}$ and $\frac{8\hat{\alpha}}{(\hat{\alpha}+1)^{2}}<\alpha$ for CH-Logit dynamic and $\hat{\gamma}<\frac{\gamma}{2}$ for CH-NTP dynamic) separate the stable region into two pieces while the total size of the region is increased.
When more travelers tend to strategically predict others’ behaviors (i.e., a larger $p^{1}$), the system becomes less stable and behaves like the over-prediction case.
Numerical experiments pertaining to the CH-Logit dynamic confirm the above theoretical results. They are omitted here for simplicity since the results and findings (such as the multi-equilibria phenomenon) are similar to the case of the CH-NTP dynamic, except that the DUE becomes the SUE.

### v2

@@ 6 CH-Logit Dynamic
While the deterministic CH-NTP dynamic captures strategic reasoning under perfect information, real-world route choices involve perceptual errors and decision uncertainty. We now replace the day-to-day operator $H[\cdot]$ with a stochastic Logit dynamic where randomness arises from quantal responses — a concept in which travelers are more likely to choose routes with lower costs yet may still select higher-cost alternatives due to inherent perception errors (Fudenberg and Levine 1998). This extension enables us to examine whether key findings about strategic prediction behavior (such as multiple equilibria and stability conditions) remain consistent across different dynamic formulations, including both deterministic and stochastic ones.
If the day-to-day operator $H[\cdot]$ is replaced by a “stochastic” Logit dynamic (Fudenberg and Levine 1998, Sandholm 2010, Xiao et al. 2019), we have the following CH-Logit dynamic:
@@
where $\Phi^{\theta}\left[\boldsymbol{z}\right]\equiv\left(d_{w}\Phi^{\theta}_{w}\left[\boldsymbol{z}_{w}\right],w\in W\right)^{T}$, $\Phi^{\theta}_{w}\left[\boldsymbol{z}_{w}\right]\equiv(d_{w}\varphi_{rw}(\boldsymbol{z}_{w}),r\in R_{w})^{T}$, $\varphi_{rw}(\boldsymbol{z}_{w})=\frac{\exp(-\theta z_{rw})}{\sum_{s\in R_{w}}{\exp(-\theta z_{sw})}}$, and $\theta$ is a dispersal parameter capturing travelers’ perception errors. For each $w\in W$, as $\theta\to 0$, route choices become equally probable across all routes in $R_{w}$. Conversely, as $\theta\to\infty$, choices become extremely concentrated on the minimum cost route of $R_{w}$. Note that for the Logit dynamic, the flow variable (i.e., the first variable) in $\boldsymbol{y}[\cdot]$ of Eq. (1) is not explicitly required.
As per Section 2.3, the predicted flow pattern of each class $k\in\{0,1,2\}$ reads:
@@
A significantly large $\hat{\theta}$ can be interpreted as higher-step travelers anticipating that lower-step travelers will strongly favor yesterday’s shortest path. This situation bears similarities to the case of a large $\hat{\gamma}$ in the CH-NTP model. In both cases, higher-step travelers are making an extreme prediction about the responsiveness of lower-step travelers to cost differences.
@@ 6.1 Mixed prediction-based stochastic equilibria
Similar to the CH-NTP dynamic, we term fixed points of the CH-Logit dynamic as mixed prediction-based stochastic equilibria (MPSE) and describe them in the following proposition.
@@
When the route cost function $\boldsymbol{c}\left(\boldsymbol{x}\right)$ is continuous, the dynamical system in Eqs. (26)-(29) admits at least one MPSE (i.e., one fixed point). Moreover, a vector $\boldsymbol{x}^{\circ}\equiv(\boldsymbol{x}^{k,\circ},k=0,1,2)^{T}$ is an MPSE if $\forall r\in R_{w},w\in W,k\in K$,
@@
where $\boldsymbol{\pi}^{k,\circ}\equiv(\pi_{sw}^{k},s\in R_{w},w\in W)^{T}$ is a function of the aggregate flow pattern $\tilde{\boldsymbol{x}}^{\circ}=\sum_{k}{\boldsymbol{x}^{k,\circ}}$, dictated by Eqs. (27)-(29).
@@
The existence of a fixed point is guaranteed by the continuity of both the closed-form Logit operator and the route cost function. To find the fixed points of the dynamical system, we set the RHS of Eq. (26) to zero and rearrange terms, yielding Eq. (30). ∎
@@
$x_{rw}^{k,\star}=p^{k}d_{w}\frac{\exp\left(-\theta c_{rw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}{\sum_{s\in R_{w}}{\exp\left(-\theta c_{sw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}}.$ (31)
@@
When $\hat{\theta}=\theta$, a $|K|(=3)$-class SUE in Definition 2 parameterized by $\theta$ is an MPSE of the dynamical system with Eqs. (26)-(29), but not vice versa.
@@
At the SUE, we first observe from Eq. (27) that $\boldsymbol{\pi}^{0,\star}=\tilde{\boldsymbol{x}}^{\star}$. Thus, for each $r\in R_{w},w\in W$, the 1-step traveler’s predicted flow Eq. (28) becomes:
$\pi^{1,\star}_{rw}=\hat{\alpha}q_{1}^{0}d_{w}\frac{\exp\left(-\theta c_{rw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}{\sum_{s\in R_{w}}{\exp\left(-\theta c_{sw}^{k}(\tilde{\boldsymbol{x}}^{\star})\right)}}+(1-\hat{\alpha})q_{1}^{0}\tilde{x}_{rw}^{\star}=\hat{\alpha}\tilde{x}_{rw}^{\star}+(1-\hat{\alpha})\tilde{x}_{rw}^{\star}=\tilde{x}_{rw}^{\star},$ (32)
where the second equality follows directly from Definition 2.
Substituting $\boldsymbol{\pi}^{0,\star}=\tilde{\boldsymbol{x}}^{\star}$ and $\boldsymbol{\pi}^{1,\star}=\tilde{\boldsymbol{x}}^{\star}$ into Eq. (29), we derive that for each $r\in R_{w},w\in W$:
@@
Hence, $\boldsymbol{\pi}^{0,\star}$, $\boldsymbol{\pi}^{1,\star}$ and $\boldsymbol{\pi}^{2,\star}$ all equal to $\tilde{\boldsymbol{x}}^{\star}$. By substituting these three predicted costs into Eq. (26) and applying Definition 2, Eq. (26) equals 0. Therefore, the SUE constitutes an MPSE.
Similar to the CH-NTP dynamic, there exist some MPSE that are not SUE when $|K|\geq 2$. We can easily construct a counter-example demonstrating their existence (similar to Figures 1(b)-1(d)) to negate the necessity. This example is omitted to conserve space. ∎
@@ 6.2 Local stability
Due to multiple equilibria, this section examines the CH-Logit dynamic’s local stability by first deriving its Jacobian at the MPSE and then analyzing stability near the SUE.
@@ 6.2.1 Jacobian matrix
The Jacobian matrix formula for an individual OD pair $w$ follows from Gao and Pavel (2017), with a negative sign applied to account for our cost-based formulation rather than their utility-based approach, yielding $\Upsilon^{\theta}_{w}[\boldsymbol{c}_{w}]$. The block-diagonal structure results from route independence across OD pairs, and $\Upsilon^{\theta}[\boldsymbol{c}]$ is NSD since each diagonal block is NSD. ∎
@@
The derivation can be found in Appendix I. ∎
@@ 6.2.2 Stability near the SUE
When $\theta\neq\hat{\theta}$, the block submatrices in Eq. (35) do not commute with each other. This requires direct calculation of the inverse matrix of one of these submatrices to evaluate $|\lambda{I}-J\Phi|$. However, none of these inverse submatrices has a closed-form. Hence, we analytically examine the case of $\theta=\hat{\theta}$ only.
@@
Denote $\rho_{i}$ as the $i$-th eigenvalue of $\Upsilon^{\theta}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\star}\right)\right]D\left[\tilde{\boldsymbol{x}}^{\star}\right]$ at the SUE under Assumption 1 and $\rho_{i}\leq 0,\forall i$. Define function $\psi(\rho_{i};\alpha,\allowbreak\hat{\alpha},p^{0},p^{1},p^{2})\equiv\alpha\hat{\alpha}^{2}p^{2}\frac{p^{1}}{p^{0}+p^{1}}\rho_{i}^{3}+(\alpha\hat{\alpha}-\alpha\hat{\alpha}p^{0}-\alpha\hat{\alpha}^{2}p^{2}\frac{p^{1}}{p^{0}+p^{1}})\rho_{i}^{2}+(\alpha\hat{\alpha}p^{0}-\alpha\hat{\alpha}+\alpha)\rho_{i}+(1-\alpha)$ where $p^{0}+p^{1}+p^{2}=1$. When $\hat{\theta}=\theta$ and $|K|\leq 3$, the SUE is locally asymptotically stable when $-1<\psi(\rho_{i};\alpha,\hat{\alpha},p^{0},p^{1},p^{2})<1,\forall i$. For $|K|=2$, we have $p^{2}=0$. For $|K|=1$, we have $p^{1}=p^{2}=0$, $p^{0}=1$ and $\hat{\alpha}=\alpha$.
@@ 6.2.2.1   Perfect prediction
Figure 11: Stable and unstable regions of the CH-Logit dynamic with $\hat{\alpha}=\alpha$ w.r.t. different $\rho_{i}$.
@@
    When $\rho_{i}\geq-1$, the system maintains stability regardless of parameter values. This occurs because $\psi(\rho_{i}=-1;\alpha,\hat{\alpha}=\alpha,p^{0},p^{1},p^{2})=2\alpha^{2}-2(p^{0}+1)\alpha+1=(\sqrt{2}\alpha-\frac{\sqrt{2}}{2}(1+p^{0}))^{2}-\frac{1}{2}(1+p^{0})^{2}+1$, which exceeds $-1$ due to constraints $p^{0}\leq 1$ and $\alpha<1$. Further note that when $\rho_{i}\in[-1,0]$, the function monotonically increases in $\rho_{i}$. Hence, if $\rho_{i}\geq-1$, the cubic function is always greater than -1 but less than $1-\alpha$;
@@
    Given a $\rho_{i}$ in $(-\infty,-1)$, the stable region shrinks as $\alpha$ increases. Moreover, the stable region splits into two pieces when $\alpha$ exceeds a certain threshold (approximately 0.6);
@@
For case (i) of $|K|=2$ in Proposition 9, when we solve $11-\alpha+\frac{\alpha(1-\hat{\alpha}+\hat{\alpha}p^{0})^{2}}{4\hat{\alpha}(p^{0}-1)}>-$ 1 with $\hat{\alpha}\geq\alpha$, we obtain $p^{0}\leq 2\sqrt{2}\sqrt{-\frac{\alpha-2}{\alpha^{2}\hat{\alpha}^{2}}}+\frac{\alpha\hat{\alpha}+\alpha-4}{\alpha\hat{\alpha}}$. Denote the RHS of this inequality as $h(\alpha,\hat{\alpha})$. Since $h(\alpha,\hat{\alpha})$ increases monotonically with $\hat{\alpha}$ and $h(\alpha,\hat{\alpha}=\alpha)\geq 2(\sqrt{2}-1)$, we have $h(\alpha,\hat{\alpha})\geq 2(\sqrt{2}-1)\approx 0.828$. The stability condition for $|K|=1$ requires all the $\rho_{i}>\frac{\alpha-2}{\alpha}$ while the condition for $|K|=2$ needs all the $\rho_{i}>\frac{1}{(p^{0}-1)\hat{\alpha}}$. Comparison of these two thresholds yields: $\frac{1}{(p^{0}-1)\hat{\alpha}}-\frac{\alpha-2}{\alpha}<0$ when $p^{0}>\frac{\alpha\hat{\alpha}+\alpha-2\hat{\alpha}}{\alpha\hat{\alpha}-2\hat{\alpha}}$ and $\frac{1}{(p^{0}-1)\hat{\alpha}}-\frac{\alpha-2}{\alpha}>0$ when $p^{0}<\frac{\alpha\hat{\alpha}+\alpha-2\hat{\alpha}}{\alpha\hat{\alpha}-2\hat{\alpha}}$. Denote $\frac{\alpha\hat{\alpha}+\alpha-2\hat{\alpha}}{\alpha\hat{\alpha}-2\hat{\alpha}}$ as $g(\alpha,\hat{\alpha})$. Summarizing the above, the stable region expands when $g(\alpha,\hat{\alpha})<p^{0}<h(\alpha,\hat{\alpha})$ and shrinks when $0<p^{0}<g(\alpha,\hat{\alpha})$.
For case (ii) of $|K|=2$ in Proposition 9, $p^{0}>h(\alpha,\hat{\alpha})$. Combined size of the two separated stable regions over $\rho_{i}$ is $f_{0}(\alpha,\hat{\alpha},p^{0})-\frac{1}{(p-1)\hat{\alpha}}+0-f_{1}(\alpha,\hat{\alpha},p^{0})=\frac{1}{\hat{\alpha}(1-p^{0})}\cdot\left(1-\sqrt{\frac{8\hat{\alpha}(p^{0}-1)+\alpha(1+\hat{\alpha}-p^{0}\hat{\alpha})^{2}}{\alpha}}\right)$, which is always less than $\frac{2-\alpha}{\alpha}$ when $p^{0}>h(\alpha,\hat{\alpha})$. Hence, the total size of the two separated stable regions is expanded.
Figure 12 summarizes the above discussion by showing how the stable region changes when $|K|$ shifts from 1 to 2 under different values of $p^{0}$, $\alpha$, $\hat{\alpha}$. The parametric space of $p^{0}$ and $\alpha$ with a shrinking stable region enlarges as over-prediction severity increases (i.e., $\frac{\hat{\alpha}}{\alpha}$ becomes larger).
Figure 12: Stable region alterations when $|K|$ changes from 1 to 2 under different $p^{0}$, $\alpha$, $\hat{\alpha}$, and $\hat{\alpha}\geq\alpha$.
@@
1.When $\hat{\alpha}\geq\frac{\alpha}{2-\alpha}$, solving $\frac{1}{(p^{0}-1)\hat{\alpha}}<\frac{\alpha-2}{\alpha}$ yields $p^{0}>g(\alpha,\hat{\alpha})$ and solving $\frac{1}{(p^{0}-1)\hat{\alpha}}>\frac{\alpha-2}{\alpha}$ obtains $p^{0}<g(\alpha,\hat{\alpha})$. Hence, the region is expanded when $g(\alpha,\hat{\alpha})<p^{0}<h(\alpha,\hat{\alpha})$ and shrunk when $p^{0}<g(\alpha,\hat{\alpha})$. It can be verified that $g(\alpha,\hat{\alpha})$ is monotonically increasing in $\hat{\alpha}$ and that $h(\alpha,\hat{\alpha})-g(\alpha,\hat{\alpha})$ is monotonically decreasing in $\hat{\alpha}$ when $10<\alpha,\hat{\alpha}<$ 1. Hence, the parametric space of $p^{0}$, $\alpha$, and $\hat{\alpha}$ ($<\alpha$) rendering the stable region shrunk is smaller than the perfect prediction case, while the parametric space leading to an expanded stable region is larger than the perfect prediction case;
2.When $\hat{\alpha}<\frac{\alpha}{2-\alpha}$, $\frac{1}{(p^{0}-1)\hat{\alpha}}<\frac{\alpha-2}{\alpha}$ always holds and thus the expanded region is simply $p^{0}<h(\alpha,\hat{\alpha})$. When prediction behaviors are considered, no stable region of $|K|=1$ is shrunk.
For case (ii) where $f^{\ast}(\alpha,\hat{\alpha},p^{0})<-1$, $p^{0}>h(\alpha,\hat{\alpha})$, we verify that $f_{0}(\alpha,\hat{\alpha},p^{0})-\frac{1}{(p-1)\hat{\alpha}}+0-f_{1}(\alpha,\hat{\alpha},p^{0})>\frac{2-\alpha}{\alpha}$. Hence, when $p^{0}>h(\alpha,\hat{\alpha})$, the stable region separates and increases in combined size. Since $h(\alpha,\hat{\alpha})$ is monotonically increasing in $\hat{\alpha}$, the parametric space rendering the total size of the stable region greater is larger than the perfect prediction case.
@@
Figure 13: Stable region alterations when $|K|$ changes from 1 to 2 under different $p^{0}$, $\alpha$, $\hat{\alpha}$ and $3-2\sqrt{2}<\hat{\alpha}\leq\alpha$.
@@
This scenario depends on the comparison between $\frac{8\hat{\alpha}}{(1+\hat{\alpha})^{2}}$ and $\alpha$. When $\frac{8\hat{\alpha}}{(1+\hat{\alpha})^{2}}>\alpha$, the analysis mirrors the case of $\hat{\alpha}>3-2\sqrt{2}$; i.e., it helps stabilize the dynamic. When $\frac{8\hat{\alpha}}{(1+\hat{\alpha})^{2}}<\alpha$, $f^{\ast}(\alpha,\hat{\alpha},p^{0})<-1$ always holds and thus the stable region is separated, while the combined size is increased. The latter case is similar to the CH-NTP dynamic with $\hat{\gamma}<\frac{\gamma}{2}$.